# CaTS-Bench: Can Language Models Describe Numeric Time Series?

### v1

@@ F.1 Linguistic Metrics
$\text{F1}_{\text{DeBERTa}}=\frac{2\cdot P\cdot R}{P+R},\quad P=\frac{1}{|c|}\sum_{i\in c}\max_{j\in r}\cos(\mathbf{e}_{i},\mathbf{e}_{j}),\quad R=\frac{1}{|r|}\sum_{j\in r}\max_{i\in c}\cos(\mathbf{e}_{j},\mathbf{e}_{i})$ (1)
@@
$\text{BLEU}=\text{BP}\cdot\exp\left(\sum_{n=1}^{N}w_{n}\log p_{n}\right),\quad\text{BP}=\begin{cases}1,&\text{if }c>r\\ e^{1-r/c},&\text{if }c\leq r\end{cases}$ (2)
@@
$\text{ROUGE-L}_{\text{F1}}=\frac{(1+\beta^{2})\cdot\text{LCS}}{r+c},\quad\text{LCS}=\text{LongestCommonSubsequence}(r,c)$ (3)
@@
$\text{METEOR}=F_{\text{mean}}\cdot(1-\text{Pen}),\quad F_{\text{mean}}=\frac{10\cdot P\cdot R}{R+9P},\quad\text{Pen}=0.5\left(\frac{\text{chunks}}{\text{matches}}\right)^{3}$ (4)
@@ F.2 Tolerance Design in Numeric Metrics
Appendix G Full TSC Results
Here, we report the full evaluation results for time series captioning, using human-revisited and semi-synthetic captions as ground truth. See Tables 7 and 8 respectively.
Table 7: Evaluation of generated captions against the human-revisited ground truths. Numeric: numeric score. Mean/STD/Max/Min refer to statistical inference accuracy. Bolded and underlined scores denote first and second places.
Model DeBERTa F1 SimCSE BLEU ROUGE-L METEOR Mean STD Max Min Numeric
Proprietary
Gemini 2.0 Flash 0.6645 0.8558 0.0793 0.2475 0.2205 0.5357 - 0.9823 0.9363 0.6335
Gemini 2.5 Pro Prev. 0.6568 0.8570 0.0690 0.2363 0.2468 0.3229 - 0.9871 0.9771 0.6805
Claude 3 Haiku 0.6580 0.8525 0.0640 0.2405 0.2358 0.8333 - 0.9797 0.9339 0.6007
GPT-4o 0.6605 0.8632 0.0705 0.2328 0.2355 0.8167 - 0.9921 0.9379 0.6268
Pretrained
InternVL 2.5 (8b) 0.6418 0.8411 0.0510 0.2093 0.2153 0.7813 0.0000 0.9495 0.8186 0.5812
InternVL 2.5 (38b) 0.6640 0.8707 0.0723 0.2442 0.2550 0.8581 0.0000 0.9820 0.9297 0.6585
LLaVA v1.6 Mistral 0.6268 0.8243 0.0525 0.2145 0.2333 0.6667 0.0000 0.8714 0.7505 0.4548
LLaVA v1.6 34b 0.6388 0.8205 0.0595 0.2210 0.2317 0.4103 0.5588 0.8170 0.7274 0.5465
Phi-4 M.I. 0.6168 0.8196 0.0450 0.2355 0.1945 0.4554 0.2500 0.9227 0.8926 0.5533
Idefics 2 0.6023 0.7838 0.0235 0.1918 0.1400 0.8056 0.3854 0.8908 0.8401 0.4238
SmolVLM 0.5918 0.7552 0.0273 0.1935 0.1538 0.9050 0.5278 0.8972 0.7606 0.4308
QwenVL 0.6185 0.8205 0.0490 0.2088 0.2140 0.6563 - 0.7947 0.6776 0.4450
QwenVL PAL 0.6643 0.8643 0.0660 0.2373 0.2260 0.9730 0.8571 0.9848 0.9784 0.5638
Llama 3.2 Vision 0.6527 0.8520 0.0715 0.2390 0.2515 0.4667 - 0.9562 0.8952 0.6503
Gemma 3 12b 0.6568 0.8692 0.0713 0.2350 0.2633 0.8148 1.0000 0.9576 0.8956 0.6598
Gemma 3 27b 0.6480 0.8634 0.0650 0.2220 0.2565 0.7341 - 0.9780 0.9035 0.6407
Finetuned
InternVL 2.5 (8b) 0.6378 0.8171 0.0533 0.2148 0.2285 0.7500 0.0000 0.8302 0.7339 0.5815
LLaVA v1.6 Mistral 0.7123 0.8964 0.1335 0.3123 0.3003 0.9284 0.1746 0.9866 0.9806 0.6932
Phi-4 M.I. 0.6485 0.8501 0.0638 0.2393 0.2295 0.6250 0.2500 0.9655 0.9184 0.5860
Idefics 2 0.7108 0.8942 0.1323 0.3085 0.2975 0.9580 0.3453 0.9875 0.9665 0.6912
SmolVLM 0.6035 0.8168 0.0513 0.2275 0.2198 0.6399 0.1250 0.9139 0.7721 0.5558
QwenVL 0.7032 0.8920 0.1263 0.3020 0.2965 0.9520 0.2430 0.9731 0.9634 0.6825
Llama 3.2 Vision 0.6470 0.8518 0.0693 0.2328 0.2483 0.5876 0.5000 0.9585 0.8961 0.6273
Table 8: Evaluation of generated captions against semi-synthetic ground truths. Numeric: numeric score. Mean/STD/Max/Min refer to statistical inference accuracy. Bolded and underlined scores denote first and second places.
Model DeBERTa F1 SimCSE BLEU ROUGE-L METEOR Mean STD Max Min Numeric
Proprietary
Gemini 2.0 Flash 0.688 0.858 0.137 0.318 0.279 0.651 0.916 0.985 0.917 0.677
Gemini 2.5 Pro Prev. 0.668 0.845 0.088 0.267 0.284 0.494 0.667 0.994 0.971 0.714
Claude 3 Haiku 0.682 0.856 0.112 0.291 0.300 0.693 0.735 0.977 0.898 0.623
GPT-4o 0.681 0.865 0.112 0.284 0.296 0.700 0.778 0.990 0.921 0.644
Pretrained
InternVL 2.5 (8b) 0.659 0.794 0.081 0.247 0.260 0.610 0.920 0.949 0.794 0.589
InternVL 2.5 (38b) 0.688 0.868 0.129 0.305 0.331 0.784 0.640 0.966 0.887 0.685
LLaVA v1.6 Mistral 0.650 0.820 0.086 0.259 0.287 0.644 0.611 0.864 0.743 0.517
LLaVA v1.6 34b 0.655 0.825 0.094 0.265 0.285 0.445 0.550 0.843 0.698 0.560
Phi-4 M.I. 0.624 0.797 0.074 0.274 0.239 0.457 0.443 0.942 0.859 0.583
Idefics 2 0.604 0.698 0.040 0.226 0.162 0.616 0.368 0.903 0.806 0.455
SmolVLM 0.594 0.693 0.044 0.224 0.178 0.747 0.446 0.864 0.705 0.474
QwenVL 0.643 0.890 0.082 0.249 0.261 0.565 0.257 0.822 0.657 0.504
QwenVL PAL 0.685 0.843 0.108 0.292 0.282 0.903 0.549 0.980 0.942 0.613
Llama 3.2 Vision 0.671 0.850 0.118 0.290 0.315 0.594 0.666 0.952 0.877 0.685
Gemma 3 12b 0.676 0.867 0.097 0.279 0.317 0.653 0.578 0.957 0.879 0.673
Gemma 3 27b 0.667 0.863 0.085 0.263 0.309 0.694 0.900 0.968 0.864 0.668
Finetuned
InternVL 2.5 (8b) 0.655 0.809 0.088 0.259 0.282 0.597 0.464 0.904 0.779 0.594
LLaVA v1.6 Mistral 0.758 0.907 0.285 0.445 0.441 0.828 0.294 0.976 0.926 0.732
Phi-4 M.I. 0.662 0.821 0.010 0.285 0.279 0.645 0.641 0.965 0.877 0.607
Idefics 2 0.759 0.908 0.290 0.452 0.437 0.885 0.379 0.985 0.927 0.733
SmolVLM 0.613 0.781 0.091 0.269 0.265 0.590 0.297 0.898 0.777 0.643
QwenVL 0.643 0.790 0.082 0.249 0.260 0.565 0.257 0.822 0.657 0.504
Llama 3.2 Vision 0.667 0.844 0.111 0.283 0.310 0.502 0.619 0.955 0.867 0.668
Appendix H Quality Validation
H.1 Manual Verification of Semi-synthetic Ground Truth Captions
To assess the reliability of semi-synthetic captions used as ground truth, we manually verified statistical and trend claims in around $2.9k$ captions of the test set under a three-tier scoring system: exact (within $\pm 0.05$ of the true value), near (within $\pm 0.1$), and incorrect.
@@ H.1 Manual Verification of Semi-synthetic Ground Truth Captions
Task Feature/Type Occurences Verified Accuracy
@@
H.2 Human Study on Detectability
We conducted a blind study with $35$ participants, each reviewing 11 captions (half written by humans, half by Gemini, using the same context information). Participants labeled each as human or AI-written but achieved only $41.1\%$ accuracy, essentially random, suggesting Gemini’s captions were indistinguishable from human ones. Participation form with guidelines is in Appendix O.
H.3 Robustness of Evaluation - Paraphrasing Experiment
@@ 3.2 Vision
We systematically paraphrased a representative subset of our ground truth captions (agriculture, crime, demography, Walmart) using architecturally distinct LLMs (GPT-4o, Gemma 27B, and Llama 90B), resulting in three additional sets of ground truth captions. The paraphrasing prompt was designed to instruct the model to thoroughly alter sentence structure, syntax, and word choice while preserving all factual information, numeric values, and statistical details with absolute fidelity. The prompt used for paraphrasing is shown in N.3. We define this paraphrased caption set as CaTS-Bench-Paraphrased, which contains captions that are semantically equivalent to the original ground truth but differ only in linguistic style.
We then re-evaluated the outputs of all six representative pretrained models against CaTS-Bench-Paraphrased using our full suite of metrics. Results of this analysis are presented in the Table 10, one for each ground-truth generator’s linguistic style. Values represent the average across selected domains. To ease comparison, we also report the results obtained with our original Gemini captions as ground truth.
@@
Table 11: Spearman correlation of model rankings of Gemini as ground truth vs. different models as ground truth. A Spearman Correlation of 11 means ranking does not change at all.
DeBERTa F1 SimCSE BLEU ROUGE-L METEOR Average
Spearman Correlation 0.9714 1.0000 0.8138 0.9048 0.9429 0.9266
p-value 0.0018 0.0000 0.0557 0.0257 0.0079 0.0182
@@
H.4 Diversity Analysis
H.4.1 Content in the Captions
@@
Table 12: Coverage of statistical and trend descriptor clusters across benchmark captions. Captions consistently include diverse descriptors capturing both quantitative and temporal aspects of the data.
@@
Trend and Patterns Peaks and Valleys (peak, dip, spike, trough) 3013
@@
H.4.2 N-Gram Diversity
We measured type token ratios (TTR) across $411k$ tokens in the semi-synthetic test set. The ratios rise significantly with $n$: TTR = $0.0288$ (1 gram), $0.1288$ (2 gram), $0.2971$ (3 gram), $0.4638$ (4 gram), and $0.6050$ (5 gram), indicating that the phrases diversify rapidly as $n$ increases. Notably, we observe over $250k$ unique 5-grams out of $411k$ tokens, providing strong evidence that the captions are not overly templated.
H.4.3 Latent Similarities

### v2

@@ F.1 Linguistic Metrics
$\text{F1}_{\text{DeBERTa}}$ $=\frac{2PR}{P+R},$ (1)
$\text{where}\quad P$ $=\frac{1}{|c|}\sum_{i\in c}\max_{j\in r}\cos(\mathbf{e}_{i},\mathbf{e}_{j}),$
$R$ $=\frac{1}{|r|}\sum_{j\in r}\max_{i\in c}\cos(\mathbf{e}_{j},\mathbf{e}_{i}).$
@@
BLEU $=\text{BP}\cdot\exp\!\left(\sum_{n=1}^{N}w_{n}\log p_{n}\right),$ (2)
[…]
@@
$\text{ROUGE-L}_{\text{F1}}$ $=\frac{(1+\beta^{2})\cdot\text{LCS}}{r+c}$ (3)
@@
METEOR $=F_{\text{mean}}\cdot(1-\text{Pen}),$ (4)
$\text{where}\quad F_{\text{mean}}$ $=\frac{10\cdot P\cdot R}{R+9P},$
[…]
@@ F.2 Tolerance Design in Numeric Metrics
F.3 Attempt Analysis in Numeric Metrics
We provide a detailed diagnostic analysis of model behavior across the four numeric inference tasks: mean, minimum (Min), maximum (Max), and standard deviation (STD). We analyze each statistic independently to expose distinct reasoning behaviors, robustness patterns, and failure modes. All observations are grounded in successes, attempts, and omissions (Null) of these statistics.
Specifically, success rate is defined as the proportion of correctly reported statistical claims among attempts. To illustrate, if the model only attempted to report the time series mean once across the entire test set, and the mean is correct, the success rate would be $1$. To account for this flaw, the attempt rate is defined as the proportion of times where the model does make the statistical claim, whether it is correct or not. By multiplying the success rate by the attempt rate, we obtain the proportion of correct statistical claims over all evaluation instances. Finally, we count the statistical omissions (Null) where the model did not attempt the statistical claim whatsoever.
Table 6: Mean inference performance across domains and models.
Domain Model Success Attempt Null
Agriculture Gemini 2.0 Flash 0.9940 1.0000 0
GPT-4o 0.9940 1.0000 0
QwenVL 0.4850 0.9581 7
QwenVL PAL 0.9701 0.9701 5
Border Crossing Gemini 2.0 Flash 0.7709 0.9907 2
GPT-4o 0.8359 0.9938 1
QwenVL 0.2879 0.9876 4
QwenVL PAL 0.9659 0.9721 9
COVID Gemini 2.0 Flash 0.5167 0.9700 9
GPT-4o 0.5367 1.0000 0
QwenVL 0.1200 0.9233 23
QwenVL PAL 0.8400 0.9833 5
Crime Gemini 2.0 Flash 0.8235 1.0000 0
GPT-4o 0.9281 1.0000 0
QwenVL 0.2092 0.9673 5
QwenVL PAL 0.9804 0.9804 3
Demography Gemini 2.0 Flash 0.9917 1.0000 0
GPT-4o 1.0000 1.0000 0
QwenVL 0.6000 0.6333 44
QwenVL PAL 0.9500 0.9500 6
Diet Gemini 2.0 Flash 0.9763 0.9858 3
GPT-4o 0.9858 0.9858 3
QwenVL 0.6066 0.9005 21
QwenVL PAL 0.9242 0.9336 14
Road Injuries Gemini 2.0 Flash 0.9868 0.9868 2
GPT-4o 1.0000 1.0000 0
QwenVL 0.3816 0.9145 13
QwenVL PAL 0.9474 0.9539 7
Walmart Gemini 2.0 Flash 1.0000 1.0000 0
GPT-4o 0.9817 1.0000 0
QwenVL 0.4128 0.7431 28
QwenVL PAL 0.9908 0.9908 1
Table 7: Minimum inference performance across domains and models.
Domain Model Success Attempt Null
Agriculture Gemini 2.0 Flash 0.9641 0.9641 6
GPT-4o 0.9940 0.9940 1
QwenVL 0.3533 0.8503 25
QwenVL PAL 0.9641 0.9641 6
Border Crossing Gemini 2.0 Flash 0.9721 0.9907 2
GPT-4o 0.9876 0.9969 0
QwenVL 0.3839 0.9505 16
QwenVL PAL 0.9505 0.9876 4
COVID Gemini 2.0 Flash 0.9000 0.9600 12
GPT-4o 0.9567 0.9933 2
QwenVL 0.3533 0.8133 56
QwenVL PAL 0.9633 0.9800 6
Crime Gemini 2.0 Flash 0.9869 0.9935 1
GPT-4o 1.0000 1.0000 0
QwenVL 0.1765 0.9804 3
QwenVL PAL 0.9804 0.9869 2
Demography Gemini 2.0 Flash 0.9667 0.9667 4
GPT-4o 0.9917 0.9917 1
QwenVL 0.5083 0.5083 59
QwenVL PAL 0.9333 0.9417 7
Diet Gemini 2.0 Flash 0.9668 0.9763 5
GPT-4o 0.9763 0.9763 5
QwenVL 0.5261 0.8863 24
QwenVL PAL 0.9289 0.9526 10
Road Injuries Gemini 2.0 Flash 1.0000 1.0000 0
GPT-4o 1.0000 1.0000 0
QwenVL 0.6382 0.7829 33
QwenVL PAL 0.9605 0.9605 6
Walmart Gemini 2.0 Flash 0.9817 0.9817 2
GPT-4o 1.0000 1.0000 0
QwenVL 0.6606 0.9908 1
QwenVL PAL 0.9908 0.9908 1
Table 8: Maximum inference performance across domains and models.
Domain Model Success Attempt Null
Agriculture Gemini 2.0 Flash 0.9760 0.9760 4
GPT-4o 0.9940 0.9940 1
QwenVL 0.9042 0.9880 2
QwenVL PAL 0.9820 0.9940 1
Border Crossing Gemini 2.0 Flash 0.9783 0.9907 2
GPT-4o 0.9907 0.9969 0
QwenVL 0.7771 0.9783 7
QwenVL PAL 0.9536 0.9938 2
COVID Gemini 2.0 Flash 0.9667 0.9733 8
GPT-4o 0.9900 0.9933 2
QwenVL 0.5367 0.8933 32
QwenVL PAL 0.9800 1.0000 0
Crime Gemini 2.0 Flash 0.9935 0.9935 1
GPT-4o 1.0000 1.0000 0
QwenVL 0.7386 0.9935 1
QwenVL PAL 0.9804 1.0000 0
Demography Gemini 2.0 Flash 0.9667 0.9667 4
GPT-4o 0.9917 0.9917 1
QwenVL 0.6833 0.6917 37
QwenVL PAL 0.9333 0.9750 3
Diet Gemini 2.0 Flash 0.9810 0.9810 4
GPT-4o 0.9810 0.9810 4
QwenVL 0.8341 0.9479 11
QwenVL PAL 0.9336 0.9953 1
Road Injuries Gemini 2.0 Flash 0.9737 0.9737 4
GPT-4o 1.0000 1.0000 0
QwenVL 0.8487 0.9079 14
QwenVL PAL 0.9605 0.9934 1
Walmart Gemini 2.0 Flash 0.9725 1.0000 0
GPT-4o 1.0000 1.0000 0
QwenVL 0.2569 0.9908 1
QwenVL PAL 0.9908 1.0000 0
Table 9: Standard deviation inference performance across domains and models.
Domain Model Success Attempt Null
Agriculture Gemini 2.0 Flash 0.0000 0.9940 167
GPT-4o 0.0000 1.0000 167
QwenVL 0.0000 0.7246 167
QwenVL PAL 0.0000 0.9581 167
Border Crossing Gemini 2.0 Flash 0.4149 0.9412 18
GPT-4o 0.4551 0.9907 2
QwenVL 0.0248 0.9876 4
QwenVL PAL 0.7678 0.9659 11
COVID Gemini 2.0 Flash 0.3633 0.9700 9
GPT-4o 0.3367 1.0000 0
QwenVL 0.0100 0.9267 22
QwenVL PAL 0.8000 0.9567 13
Crime Gemini 2.0 Flash 0.4771 0.9935 1
GPT-4o 0.4771 1.0000 0
QwenVL 0.0458 0.9869 2
QwenVL PAL 0.9608 0.9804 3
Demography Gemini 2.0 Flash 0.5167 1.0000 0
GPT-4o 0.2833 0.9917 1
QwenVL 0.0417 0.8833 14
QwenVL PAL 0.0333 0.9333 8
Diet Gemini 2.0 Flash 0.0000 0.9716 211
GPT-4o 0.0000 0.9953 211
QwenVL 0.0000 0.7536 211
QwenVL PAL 0.0000 0.9336 211
Road Injuries Gemini 2.0 Flash 0.2566 0.9605 6
GPT-4o 0.1316 1.0000 0
QwenVL 0.0395 0.9342 10
QwenVL PAL 0.0197 0.9013 15
Walmart Gemini 2.0 Flash 0.0000 0.9725 109
GPT-4o 0.0000 0.9817 109
QwenVL 0.0000 0.8073 109
QwenVL PAL 0.0000 0.9908 109
Mean Inference.
Mean inference is the most stable and reliable numeric operation across all evaluated models. Strong models consistently achieve high success rates ($\geq 0.97$) with near-universal attempt rates ($\approx 1.00$) and negligible null counts across all settings. For example, Gemini2.0 Flash and GPT-4o attain success rates between $0.97$ and $1.00$ with zero or near-zero nulls throughout Table 6. Weaker models exhibit a distinct pattern: while attempt rates remain high (typically 0.900.900.980.98), success drops substantially (e.g., 0.120.120.610.61), accompanied by elevated null counts (often >20>20). This decoupling of attempt and success indicates that failures arise from imprecise aggregation rather than abstention. Importantly, mean inference performance does not correlate with time-series length. Domains with long horizons show comparable success and null rates to short-horizon settings, confirming that temporal span is not a limiting factor.
Min Inference.
[…]
Max Inference.
Maximum inference is the most robust extrema-related operation. Across nearly all models and settings, attempt rates are $\approx 1.00$, omission is minimal, and success rates are consistently high. Strong models achieve success rates of 0.970.971.001.00, while even weaker models often exceed 0.730.730.900.90 success (Table 8). Notably, high success coincides with aggressive attempt behavior: failures are not driven by abstention. For instance, QwenVL attempts maximum inference in >97>97 of cases but still exhibits reduced success in some settings, indicating semantic or conceptual errors rather than omission. Models exhibit a directional asymmetry in extrema reasoning: identifying peaks is substantially easier than identifying troughs, and high maximum scores are not artifacts of selective answering.
STD Inference.
Standard deviation inference exposes a fundamental capability gap across all evaluated models. Unlike the other statistics, models frequently attempt STD inference (attempt rates often 0.900.901.001.00), yet produce null or unscorable outputs in the majority of cases. This results in near-zero success across many settings. For example, several domains exhibit success =0.00=0.00 with null counts exceeding 100100200200 despite high attempt rates (Table 9). This failure pattern is not driven by abstention, nor by sequence length. Even short-horizon settings show complete failure. PAL increases attempt rates but rarely improves success, demonstrating that the issue is not instruction following but the absence of internalized variance modeling. STD inference requires explicit reasoning about dispersion, which current models do not reliably internalize. Unlike extrema, variance cannot be approximated through semantic heuristics.
Appendix G Quality Validation
G.1 Manual Verification of Synthetic Captions
To assess the reliability of synthetic captions used as ground truth, we manually verified statistical and trend claims in around $2.9k$ captions of the test set under a three-tier scoring system: exact (within $\pm 0.05$ of the true value), near (within $\pm 0.1$), and incorrect.
@@ G.1 Manual Verification of Synthetic Captions
Task Feature/Type Occurences Acc.
@@
G.2 Human Study on Detectability
We conducted a blind study with $35$ participants, each reviewing 11 captions (half written by humans, half by Gemini, using the same context information). Participants labeled each as human or AI-written but achieved only $41.1\%$ accuracy, essentially random, suggesting Gemini’s captions were indistinguishable from human ones. Participation form with guidelines is in Appendix M.
G.3 Robustness of Evaluation - Paraphrasing Experiment
@@
We systematically paraphrased a representative subset of our ground truth captions (agriculture, crime, demography, Walmart) using architecturally distinct LLMs (GPT-4o, Gemma 27B, and Llama 90B), resulting in three additional sets of ground truth captions. The paraphrasing prompt was designed to instruct the model to thoroughly alter sentence structure, syntax, and word choice while preserving all factual information, numeric values, and statistical details with absolute fidelity. The prompt used for paraphrasing is shown in L.3. We define this paraphrased caption set as CaTS-Bench-Paraphrased, which contains captions that are semantically equivalent to the original ground truth but differ only in linguistic style.
We then re-evaluated the outputs of all six representative pretrained models against CaTS-Bench-Paraphrased using full suite of metrics. Results of this analysis are presented in the Table 11, one for each ground-truth generator’s linguistic style. Values represent the average across select domains. We also report the results obtained with our original Gemini captions as ground truth.
@@
Table 12: Spearman correlation between Gemini-based rankings and rankings induced by different metrics. A value of 11 indicates identical rankings.
Metric Correlation $p$-value
DeBERTa F1 0.971 0.0018
SimCSE 1.000 0.0000
BLEU 0.814 0.0557
ROUGE-L 0.905 0.0257
METEOR 0.943 0.0079
Average 0.927 0.0182
@@
G.4 Diversity Analysis
G.4.1 Content in the Captions
@@
We apply the same keyword-based analysis to the HR caption set to examine the types of statistical and temporal descriptors used. Across domains, human captions draw from a wide range of descriptor clusters, including measures of central tendency, extrema, dispersion, and multiple forms of trend characterization. In addition to core statistical terms, human captions frequently include compound trend expressions, contextual qualifiers, and multi-stage temporal descriptions that span increases, decreases, stabilization, and recovery phases within a single caption.
Coverage extends beyond surface-level descriptors to include higher-order patterns such as structural transitions, regime changes, and shape-based trends, which appear with greater regularity in the human set. While all human captions contain at least one statistical or trend-related element, many incorporate several distinct clusters simultaneously, reflecting a richer and more varied composition of quantitative and temporal content.
Table 13: Coverage of statistical and temporal descriptor clusters across human captions. For each cluster, representative surface-form keywords are shown (non-exhaustive); keyword lists are illustrative
Category Cluster (Representative Keywords) Captions
Statistical Central Tendency (average, mean, median, mode, count, etc.) 3184
Maximum Values (max, maximum, peak, highest value, record high, etc.) 1487
Minimum Values (min, minimum, dip, trough, lowest value, etc.) 1530
Dispersion (standard deviation, deviation, variance, std, iqr, etc.) 684
Variability (variability, variation, fluctuate, oscillation, noise, etc.) 665
Range / Spread (range, spread, difference, gap, extent, etc.) 341
Extremes (peak, spike, trough, extreme, record high, etc.) 2966
Trend and Patterns Increasing Trends (increase, rising, growing, upward, gain, etc.) 1207
Decreasing Trends (decrease, declining, falling, drop, reduction, etc.) 1308
Stability / Flat Trends (stable, steady, flat, constant, plateau, etc.) 1082
Sudden Changes (spike, jump, shock, abrupt, rapid increase, etc.) 219
Volatility (fluctuate, volatile, oscillate, erratic, instability, etc.) 1210
Comparative Relationships (higher than, lower than, exceeds, below, gap, etc.) 1433
Structural Changes (inflection, shift, transition, regime change, pivot, etc.) 190
Recovery / Correction (recovery, rebound, normalize, correction, bounce back, etc.) 362
Cyclical Patterns (seasonal, cyclical, periodic, recurring, repeats, etc.) 135
Shape-based Patterns (v-shaped, u-shaped, w-shaped, rise-and-fall, hill-shaped, etc.) 75
Table 14: Coverage of statistical and trend descriptor clusters across synthetic captions. Captions consistently include diverse descriptors capturing both quantitative and temporal aspects of the data.
@@
Trend &
Patterns
Peaks and Valleys (peak, dip, spike, trough) 3013
@@
G.4.2 N-Gram Diversity
We analyze n-gram diversity using type–token ratio (TTR) to assess whether the generated captions exhibit repetitive or templated phrasing. For the synthetic corpus, computed over approximately 411411k tokens, TTR increases steadily with n-gram order, rising from $0.0288$ for unigrams to $0.6050$ for 5-grams. This monotonic growth indicates rapid diversification at the phrase level rather than reliance on fixed templates. In particular, the presence of over 250250k unique 5-grams out of 411411k tokens suggests substantial higher-order lexical variability in the synthetic captions.
We perform the same analysis on the human-revisited gold captions to provide a reference point. As shown in Table 15, HR captions exhibit consistently higher TTR across all n-gram orders, reflecting the greater compositional variability expected from human-authored text. Importantly, the synthetic captions follow a similar upward trajectory with increasing n, differing primarily in magnitude rather than qualitative behavior. This pattern suggests that while synthetic captions are less diverse than their human-revisited counterparts, they remain far from being rigidly templated and capture a broad range of phrase structures.
Table 15: Type–token ratio (TTR) for synthetic and human-revisited captions across n-gram orders. TTR increases monotonically with n for both corpora, indicating growing phrase-level diversity. Human-revisited captions exhibit higher TTR across all n, while synthetic captions remain highly non-templated at higher n-grams.
N-gram Synthetic TTR Human TTR
1-gram 0.0288 0.0472
2-gram 0.1288 0.2320
3-gram 0.2971 0.5011
4-gram 0.4638 0.7066
5-gram 0.6050 0.8344
G.4.3 Latent Similarities